You deposited BDT 120,000 in a bank account 7 years ago. The bank had been giving you 12% interest

compounded quarterly for the last 7 years. Today the bank revised the interest rate and the new interest rate

will be 6% interest compounded monthly from today onwards. How much money will you have in your

account after 4 years from now?


1
Expert's answer
2021-08-11T07:23:18-0400

Amount deposited 7 years ago = 120,000

Interest rate 7 years ago = 12% 

That is compounded quarterly which means four times in a year.

 

Computation of value after 7 years will be:

"Value \space after\space 7\space years = Value×(1+Rate)^{Time} \\ = 120,000 ×(1+\frac{0.12}{4})^{7×4} \\ =120,000 ×(1+0.03)^{28} \\ =120,000 ×2.2879276757 \\ =274,551.321088"

"Value\space of\space deposit = \$274,551.32 ×(1+\frac{0.06}{12})^{4×12}\\ = \$274,551.32 ×(1+0.005)^{48}\\ =\$274,551.32 × 1.2704891610 \\ = \$348,814.47619824252"

New interest now = 6% 

Compounded on a monthly basis 

Value of deposit after 4 years as follows:


The value of the deposit will be $348,814.48.